A trolley rolls down a slope at a constant acceleration.

It travels 8.0 m in 4.0 s.

(a) Plot the velocity-time graph for the motion of the

trolley

 

(b) Find the acceleration of the trolley


1
Expert's answer
2022-05-19T05:43:21-0400

The acceleration of the trolley

"a=2d/t^2=2*8.0/4.0^2=1.0\:\rm m/s^2"

The final speed of a trolley

"v=at=1.0*4.0=4.0\:\rm m/s"